# No Bells, Just Whistles: Sports Field Registration by Leveraging Geometric Properties

### v1

@@ 2 Related Work
Sports field registration is a critical component of most sports applications in computer vision, whose common approaches intend to estimate homography matrices in team sports. Homography estimation has traditionally relied on the corresponding features, or keypoints, identification between images and the court field model. These keypoints, usually obtained by exploiting geometric primitives such as lines and/or circles, are subsequently used to estimate the mapping between the images by using RANSAC [14] algorithm with DLT [18] or non-linear optimization through the minimisation of a chosen loss function. More recent approaches either directly predict an initial homography matrix. Alternatively, they seek the optimal matching homography within a reference database that includes synthetic images with known homography matrices or camera parameters.
@@ 3.1 Modelling the Soccer Field
A soccer field is composed of lines and circle segments, representing all field markings, goal posts, and crossbars. Our approach, like keypoint-based methods [9, 25, 7, 22, 10, 26], relies on the lines painted on the ground, their intersections, and the corners they define, due to its known position on the world coordinate system. We follow the segment definitions of Cioppa et al. [8], and set them as starting points to hierarchically compute our pre-defined keypoints from court geometric properties.
@@ 3.1.1 Keypoint Generation
  • Line-Line intersections. This set of points ($\mathcal{K}p$) includes the intersections of boundary lines or the penalty area markings. Considering the 23 lines depicted in [8], including goal posts and crossbars, up to 30 points can be included.
  • Extended Line-Line intersections. This set ($\mathcal{K}e$) addresses the intersections of extended lines that represent non-adjacent segments of the soccer field. To obtain that, the lines are extended by exploiting their equations beyond their original boundaries.
  • Line-Ellipse intersections. This set ($\mathcal{K}p_{1}$) is to consider the intersections between the field lines and the circles or semi-circles present on the court. Given the distortions introduced by the camera perspective, conics on the field are considered ellipses for equation computation. The parameters of these ellipses are fitted using the least squares method [17]. Intersection points were analytically derived using ellipse and line formulas.
  • Ellipse tangent points. Augmentation of available points is achieved through the utilization of tangent points on tangent lines, extending from a specified point to the previously defined ellipses. These tangent points (denoted by $\mathcal{K}p_{2}$) were analytically determined by employing an ellipse equation, incorporating the known coordinates of an external point.
  • Additional points. Once the previous points are inferred, and the corresponding homography, an additional set ($\mathcal{K}p_{3}$) of 9 points is integrated along the central pitch axis, encompassing the pitch center and penalty points. Additionally, 4 points are strategically placed to designate quarter turns along the central circle. Furthermore, the homography facilitates the inclusion of other points that are initially missing, addressing situations such as unannotated lines.
@@ 3.1.2 Keypoint Disambiguation
Due to the multi-view nature of the SoccerNet dataset [11] and, considering, for instance, one of the soccer field’s semi-circles, ambiguity appears in its respective $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ keypoints candidates, as shown in Fig. 2-bottom. To avoid that, we define two different strategies to handle that disambiguation depending on the total number of keypoints generated in the previous sets: when there are sufficient points in the $\mathcal{K}p\cup\mathcal{K}e$ set to infer a homography, i.e., four points, $\mathcal{K}p_{1}$ is computed first by choosing the candidates combination that minimizes a reprojection error. Then, we include $\mathcal{K}p_{1}$ in the homography estimation and apply the same strategy to $\mathcal{K}p_{2}$. Otherwise, we perform a grid-search involving both $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ candidates when $\mathcal{K}p\cup\mathcal{K}p_{e}\cup\mathcal{K}p_{1}^{*}\cup\mathcal{K}p_{2}^{*}\geq 4$, where * denotes a possible candidate combination. The grid-search iterates over all keypoints candidates in a set-wise manner to avoid unfeasible combinations and keeps the one with minimum reprojection error. It is worth noting that when none of the strategies can be applied, the current frame is deemed invalid for calibration purposes. Moreover, once we define the new sets, two additional constraints are applied in the case homography estimation or ellipse fitting is not accurate enough. Initially, we manually establish a reprojection error threshold to validate points. Subsequently, through iteration over combinations of keypoints, we construct vectors and ensure that the cross-products maintain consistent signs in both world and image coordinates. This final step is essential in cases where two distinct combinations yield valid top- and bottom-view perspectives of the field while exhibiting identical reprojection errors. Utilizing cross-products enables us to differentiate and retain the keypoint combination corresponding to the field’s top-view. The full keypoint generation process is depicted in Fig. 1-top.
@@ 3.2 Keypoints and Lines Detection
Our approach makes use of two encoder-decoder convolutional neural networks to estimate the position of the pre-defined keypoints and the soccer field lines depicted in [8] excluding conics, giving the last one an auxiliary role to enhance keypoint completeness. During inference, the former produces heatmaps for each pre-defined keypoint with a single Gaussian peak with $2px$ sigma positioned in the keypoint location, accompanied by an additional target channel. This additional channel reflects the inverse of the maximum value among the other target feature maps, ensuring that the resultant target tensor behaves as a probability distribution function at every spatial point. Meanwhile, the latter network produces heatmaps for each visible soccer field line within the frame, assigning two Gaussian peaks at the line extremities’ locations. Additionally, we introduce an extra channel, known as the boundary channel, to our heatmap following the approach outlined in [33]. This augmentation aims to enhance the efficient capture of global information regarding the soccer field and improve extremities detection, particularly near image borders. We effectively extract the positions of keypoints and line extremities from the generated heatmaps by employing a max pooling operation, drawing inspiration from the methodology proposed in [38]. This process is summarized on Fig. 1-bottom.
@@ 3.2.1 Architecture
The keypoint and line extremities detection utilized a modified HRNetV2-w48 [32] as the encoder’s backbone network. HRNetv2 [32] is a new family of convolutional networks that maintains high-resolution representations through the whole process resulting in semantically richer and spatially more precise representations. To improve the spatial resolution of the predicted heatmaps, we incorporated 2×2$\times$2 × upsampling and concatenated skip-connection features from the corresponding resolution of the convolution stem to fuse the features at different scales. The final predictions exhibit half the resolution of the original image, with softmax employed as the final activation function.
@@ 3.3.1 Camera Parameters Estimation
Extrinsic and intrinsic parameters are inferred by leveraging the coordinates of 3D object points and their corresponding 2D projections using the soccer field model as a calibration rig, following [37], which consists of a closed-form solution followed by a non-linear refinement based on the maximum likelihood criterion. To calibrate the 3D soccer field model rig, we also consider two additional vertical planes containing the goal polygons, including non-planar points such as keypoints belonging to the goal posts and crossbars. This strategy enhances completeness by providing estimations when insufficient points are on the ground plane.
To account for keypoint misdetections and other complexities in camera parameter retrieval, such as frames with only one non-planar keypoint visible, the calibration process was repeated on several subsets of keypoints. These subsets were selected based on various heuristics: full-keypoints, including all keypoints sets $\mathcal{K}p$, $\mathcal{K}p_{e}$, $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$ and $\mathcal{K}p_{3}$; main-keypoints, comprising only line-line intersections from the original SoccerNet annotations [8]; and ground-plane-keypoints, which excludes non-planar keypoints. Furthermore, we applied a grid of RANSAC [14] reprojection error thresholds to each subset. The final camera calibration values were determined through a heuristic voting process, prioritizing the method yielding a lower reprojection error, with emphasis on the full-keypoints subset.
@@ 4.2 Evaluation Metrics
The quality of estimated camera parameters or homography matrices can be evaluated both in 2D image and world spaces.
@@ 4.3 Implementation Details
Due to the absence of publicly available results on the multiple-view SN23 distribution, we trained two models from scratch: Multi-view (MV) and Single-view (SV). The latter is composed almost entirely of non-replay frames, ensuring a high percentage of central camera shots. We train separate networks for the keypoints and line extremities detection tasks on the SN23-train dataset [8]. For the MV model, we train for $200$ epochs, using an initial learning rate of $1e^{-2}$ and a batch size of $2$. For the SV model, we train for $200$ epochs, using an initial learning rate of $1e^{-5}$ and a batch size of $2$. We utilize the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. ${l}_{2}$-loss is used for heatmap regression in both neural networks. Data augmentation such as random horizontal flip, color jitter, and Gaussian noise are applied to enhance model robustness and generalization. Furthermore, we fine-tune both SV networks on the WC14 and TSWC datasets. GT homographies are transformed into our proposed keypoint sets and line extremities by projecting their respective world coordinates to the field’s ground plane. The experiments are conducted on a single NVIDIA GeForce RTX 2080 Ti GPU with $12$ GB of memory, and the implementation is carried out in the PyTorch framework.
@@ 4.4.2 Homography Estimation
For the evaluation on the TSWC-test dataset, we report the statistics from the papers [7, 23, 26] in Table 3, observing how our fine-tuned model outperforms those methods in most of the metrics, once again demonstrating the effectiveness of our approach without the need for further homography refinement.

### v2

@@ 2 Related Work
Sports field registration is a critical component of most sports applications in computer vision, whose common approaches intend to estimate homography matrices in team sports. Homography estimation has traditionally relied on the corresponding features, or keypoints, identification between images and the court field model. These features, usually obtained by exploiting geometric primitives such as lines and/or circles, are subsequently used to estimate the mapping between the images by using RANSAC [16] algorithm with DLT [20] or non-linear optimization through the minimisation of a chosen loss function. More recent approaches either directly predict an initial homography matrix. Alternatively, they seek the optimal matching homography within a reference database that includes synthetic images with known homography matrices.
@@ 3.1 Modelling the Soccer Field
A soccer field is composed of lines and circle segments, representing all field markings, goal posts, and crossbars. Our approach, like keypoint-based methods [32, 9, 27, 7, 24, 10, 28, 15], relies on the lines painted on the ground, their intersections, and the corners they define, due to its known position on the world coordinate system. We follow the segment definitions of Cioppa et al. [8], and set them as starting points to hierarchically compute our pre-defined keypoints from court geometric properties.
@@ 3.1.1 Keypoint Generation
  • Line-Line intersections. Following [32, 11, 9, 10, 28, 15], this set of points ($\mathcal{K}p$) includes the intersections of boundary lines or the penalty area markings. Considering the 23 lines depicted in [8], including goal posts and crossbars, up to 30 points can be included.
  • Extended Line-Line intersections. Following [11], this set ($\mathcal{K}e$) addresses the intersections of extended lines that represent non-adjacent segments of the soccer field. To obtain that, the lines are extended by exploiting their equations beyond their original boundaries.
  • Line-Ellipse intersections. Following [32, 11, 9, 10, 28, 15], this set ($\mathcal{K}p_{1}$) is to consider the intersections between the field lines and the circles or semi-circles present on the court. Given the distortions introduced by the camera perspective, conics on the field are considered ellipses for equation computation. The parameters of these ellipses are fitted using the least squares method [19]. Intersection points were analytically derived using ellipse and line formulas.
  • Ellipse tangent points. Following [18, 15], augmentation of available points is achieved through the utilization of tangent points on tangent lines, extending from a specified point to the previously defined ellipses. These tangent points (denoted by $\mathcal{K}p_{2}$) were analytically determined by employing an ellipse equation, incorporating the known coordinates of an external point.
  • Additional points. Following [7, 10, 24, 27, 28, 15], once the previous points and the corresponding homography are inferred, an additional set ($\mathcal{K}p_{3}$) of 9 points is integrated along the central pitch axis, encompassing the pitch center and penalty points. Additionally, 4 points are strategically placed to designate quarter turns along the central circle. Furthermore, the homography facilitates the inclusion of other points that are initially missing, addressing situations such as unannotated lines.
@@ 3.1.2 Keypoint Disambiguation
Due to the multi-view nature of the SoccerNet dataset [12] and, considering, for instance, one of the soccer field’s semi-circles, ambiguity appears in its respective $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ keypoints candidates, as shown in Fig. 2-bottom. To avoid that, we define two different strategies to handle that disambiguation depending on the total number of keypoints generated in the previous sets: when there are sufficient points in the $\mathcal{K}p\cup\mathcal{K}e$ set to infer a homography, i.e., four points, $\mathcal{K}p_{1}$ is computed first by choosing the candidates combination that minimizes a reprojection error. Then, we include $\mathcal{K}p_{1}$ in the homography estimation and apply the same strategy to $\mathcal{K}p_{2}$. Otherwise, we perform a grid-search involving both $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ candidates when $\mathcal{K}p\cup\mathcal{K}p_{e}\cup\mathcal{K}p_{1}^{*}\cup\mathcal{K}p_{2}^{*}\geq 4$, where * denotes a possible candidate combination. The grid-search iterates over all keypoints candidates in a set-wise manner to avoid unfeasible combinations and keeps the one with minimum reprojection error. It is worth noting that no more keypoint sets beyond $\mathcal{K}p$ are computed if neither of the above stated conditions are met. Moreover, once we define the new sets, two additional constraints are applied in the case homography estimation or ellipse fitting is not accurate enough. Initially, we manually establish a reprojection error threshold to validate points. Subsequently, through iteration over combinations of keypoints, we construct vectors and ensure that the cross-products maintain consistent signs in both world and image coordinates. This final step is essential in cases where two distinct combinations yield valid top- and bottom-view perspectives of the field while exhibiting identical reprojection errors. Utilizing cross-products enables us to differentiate and retain the keypoint combination corresponding to the field’s top-view. The full keypoint generation process is depicted in Fig. 1-top.
@@ 3.2 Keypoints and Lines Detection
Our approach is built upon Falaleev et al. [15] solution, which makes use of two encoder-decoder convolutional neural networks to estimate the position of the pre-defined keypoints and the soccer field lines depicted in [8] excluding conics, giving the last one an auxiliary role to enhance keypoint completeness. During inference, the former produces heatmaps for each pre-defined keypoint with a single Gaussian peak with $2px$ sigma positioned in the keypoint location, accompanied by an additional target channel. This additional channel reflects the inverse of the maximum value among the other target feature maps, ensuring that the resultant target tensor behaves as a probability distribution function at every spatial point. Meanwhile, the latter network produces heatmaps for each visible soccer field line within the frame, assigning two Gaussian peaks at the line extremities’ locations. Additionally, we introduce an extra channel, known as the boundary channel, to our heatmap following the approach outlined in [35]. This augmentation aims to enhance the efficient capture of global information regarding the soccer field and improve extremities detection, particularly near image borders. We effectively extract the positions of keypoints and line extremities from the generated heatmaps by employing a max pooling operation, drawing inspiration from the methodology proposed in [40]. This process is summarized on Fig. 1-bottom.
@@ 3.2.1 Architecture
Following [15], the keypoint and line extremities detection utilized a modified HRNetV2-w48 [34] as the encoder’s backbone network. HRNetv2 [34] is a new family of convolutional networks that maintains high-resolution representations through the whole process resulting in semantically richer and spatially more precise representations. To improve the spatial resolution of the predicted heatmaps, we incorporated 2×2$\times$2 × upsampling and concatenated skip-connection features from the corresponding resolution of the convolution stem to fuse the features at different scales. Additionally, a more lightweight version of the backbone is achieved by reducing dense representation layer size in the network’s last stages. The final predictions exhibit half the resolution of the original image, with softmax and sigmoid employed as the final activation functions for keypoints and lines, respectively.
@@ 3.3.1 Camera Parameters Estimation
Extrinsic and intrinsic parameters are inferred by leveraging the coordinates of 3D object points and their corresponding 2D projections using the soccer field model as a calibration rig, following [39], which consists of a closed-form solution followed by a non-linear refinement based on the maximum likelihood criterion. We follow  [15] to calibrate the 3D soccer field model rig: we consider two additional vertical planes containing the goal polygons, including non-planar points such as keypoints belonging to the goal posts and crossbars. Additionally, we extend this strategy to enhance completeness by providing estimations when insufficient points are on the ground plane, calibrating over the vertical planes, and subsequently transforming camera parameters to the ground-plane coordinate system.
To account for keypoint misdetections and other complexities in camera parameter retrieval, such as frames with only one non-planar keypoint visible, the calibration process was repeated on several subsets of keypoints. Similar to  [15], these subsets were selected based on various heuristics: full-keypoints, including all keypoints sets $\mathcal{K}p$, $\mathcal{K}p_{e}$, $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$ and $\mathcal{K}p_{3}$; main-keypoints, comprising only line-line intersections from the original SoccerNet annotations [8]; and ground-plane-keypoints, which excludes non-planar keypoints. Furthermore, we applied a grid of RANSAC [16] reprojection error thresholds to each subset. The final camera calibration values were determined through a heuristic voting process, prioritizing the method yielding a lower reprojection error, with emphasis on the full-keypoints subset.
@@ 4.2 Evaluation Metrics
The quality of estimated camera parameters or homography matrices can be evaluated both in 2D image and world space.
@@ 4.3 Implementation Details
Due to the absence of publicly available results on the multiple-view SN23 distribution, we trained two models from scratch: Multi-view (MV) and Single-view (SV). The latter is composed almost entirely of non-replay frames, ensuring a high percentage of central camera shots. We train separate networks for the keypoints and line extremities detection tasks on the SN23-train dataset [8]. For the MV model, we train for $200$ epochs, using an initial learning rate of $1e^{-2}$ and a batch size of $2$. For the SV model, we train for $200$ epochs, using an initial learning rate of $1e^{-5}$ and a batch size of $2$. We utilize the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. ${l}_{2}$-loss is used for heatmap regression in both neural networks. Data augmentation such as random horizontal flip, color jitter, and Gaussian noise are applied to enhance model robustness and generalization. Furthermore, we fine-tune both SV networks on the WC14 and TSWC datasets. GT homographies are transformed into our proposed keypoint sets and line extremities by projecting their respective world coordinates to the field’s ground plane. The experiments are conducted on a single NVIDIA GeForce RTX 2080 Ti GPU with $12$ GB of memory.
@@ 4.4.2 Homography Estimation
For the evaluation on the TSWC-test dataset, we report the statistics from the papers [27, 7, 25, 28] in Table 3, observing how our fine-tuned model outperforms those methods in most of the metrics, demonstrating the effectiveness of our approach without the need for further refinement.